Interest is compounded annually unless stated otherwise. Payments are at the end of the year unless stated otherwise. All bonds have a face value of $1000.

4. Goofy Gadgets makes toys. Their EPS is $4.00 per share. Estimate the share price using the price-earnings ratio (P-E) for a comparable firm, Hasbro, of 20.9.
Answer #1

Price = P/E*EPS = 20.9*4=83.6